OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10? negative five halves five halves two fifths negative two fifths none of the above

OpenStudy (anonymous):

first write this equation in terms of x.

OpenStudy (ybarrap):

$$ 2x - 5y = -10\\ -5y=-2x-10\\ y=\dfrac{-2x-10}{-5}\\ ~~={-2\over-5}x-{10 \over -5}\\ ~~={2\over5}x-(-2)\\ ~~={2\over5}x+2\\ $$ The slope, sitting right there next to the x, should popping right out! Let me know if you have an questions.